OpenStudy (anonymous):

Hi, The integral of Sin x / Cos x ?.. Is it .. 1\2 ln cosx? .. or is it just 1\2 cosx

OpenStudy (anonymous):

you can take the following : u = cos x du = -sinx dx and you'll end up with this form: \[\int\limits_{}^{}1/u du = \ln \left| u \right| + c = \ln \left| \cos x \right| + c\] did you get this? :)

OpenStudy (anonymous):

Ahh.. Ln (cosx) .. what the did i get the 1\2 from... hehe.. Thank u sstarica.

OpenStudy (anonymous):

sorry it's:\[= - \ln |cosx| + c\] I forgot about the (-) my bad ^^"

OpenStudy (anonymous):

lol, no problem ^_^